A desk is on sale for $220.80 , which is 76% less than the regular price.
What is the regular price?






Answer by Maths68(1474)

Let
Regular Price = x
Discount = 76% of x = (76/100)(x)
Discount = 76% of x = 76x/100

Regular price - Discount = Sale price
x - 76x/100=220.80
(100x-76x)/100=220.80
24x=220.80*100
24x=22080
24x/24=22080/24
x=920
Regular price is $920